### v3

# Elevating Visual Perception in Multimodal LLMs with Visual Embedding Distillation

Abstract
In recent times, the standard practice for developing MLLMs is to feed features from vision encoder(s) into the LLM and train with natural language supervision. This approach often causes models to lean towards language comprehension and undermine the rich visual perception signals present in the data, which are critical for tasks involving spatial reasoning in the domain of embodied AI and robotics. Is it possible to optimize both at the same time? In this work, we propose VisPer-LM, the first approach that infuses visual perception knowledge from expert vision encoders into the LLM’s (of an MLLM) hidden representations. We start by investigating MLLMs trained solely with natural language supervision and identify a positive correlation between the quality of visual representations within these models and their downstream performance. Given this insight, we formulate the objective during the pretraining stage in MLLMs as a coupled optimization of predictive visual embedding and next (text) token prediction. Moreover, through extensive probing, we observe improved visual representation quality due to embedding optimization, underscoring the effectiveness of our probing setup. We demonstrate that our VisPer-LM outperforms the single and multi-encoder baselines, proving our approach’s superiority over explicitly feeding the corresponding features to the LLM. In particular, VisPer-LM boosts performance by an average margin of up to 2.5% on various benchmarks, with a notable improvement of 8.7% on the Depth task in CV-Bench.
Figure 1: Different Paradigms for Incorporating Visual Information into LLMs. (a, b) Existing approaches liu2023improvedllava ; tong2024cambrian1 feed features from the visual encoder(s) into the LLM and train the model solely with natural language supervision, i.e., next token prediction (NTP) to align the embedding space of the vision encoder(s) and the LLM. (c) We propose distilling target visual information into the intermediate representations of the LLM from a set of auxiliary vision encoders ($\mathbf{E}^{\text{target}}$). We adopt a predictive embedding jepa optimization approach at selected LLM layers during training to minimize the embedding losses and the NTP loss function, resulting in a vision-centric approach to training the Multimodal Large Language Model. We only use a single base vision encoder during inference.
1 Introduction
Figure 2: Probing reveals a positive correlation between depth representation quality and performance on CV-Bench. (a) Increasing training data and using only the next-token prediction objective improves visual representation quality in the LLM, as well as downstream performance. (b) Our method, VisPer-LM, outperforms LLaVA-1.5 liu2023improvedllava in both probing and downstream tasks under the same settings.
In the last couple of years, Multimodal Large Language Model (MLLM) liu2023improvedllava ; zhu2023minigpt development has witnessed rapid growth, owing mainly to the increasing number of powerful LLMs abdin2024phi3technicalreporthighly ; dubey2024llama3herdmodels ; deepseekv2 ; qwen2 and large-scale datasets tong2024cambrian1 ; wang2023cogvlm ; chen2024far . A visual reasoning MLLM generally has three main components: vision encoder(s), a projector, and a decoder LLM. Given the presence of the decoder LLM responsible for producing the final outputs, the established practice to train MLLMs is with a next text token prediction objective gpt2 on relevant image-text datasets llava-one ; tong2024cambrian1 with supervision from the ground-truth text. Specifically, training an MLLM typically involves two primary stages liu2023llava : (i) Pre-Training (PT), training a projector to align the embeddings from the vision encoder(s) to those inside the LLM, and (ii) Instruction Fine-Tuning (IFT), training the projector and LLM on conversation data for downstream instruction following tasks. Traditional MLLMs liu2023llava usually use a pre-trained visual encoder like CLIP-ViT clip ; cherti2023reproducible ; liu2023llava to process the visual inputs.
Although effective on general visual reasoning tasks, the visual encoder’s features usually lack the fine-grained visual information required to perform tasks like spatial/depth reasoning mmvp ; diva ; jain2024vcoder , which are important visual perception abilities. Consequently, several recent works propose simply scaling the number of visual (input) encoders tong2024cambrian1 ; li2024mgm ; kar2024brave ; jain2024vcoder to incorporate auxiliary information amicable for improving the model’s visual perception abilities. However, existing works often require relatively large amounts of training data kar2024brave ; tong2024cambrian1 and compute nvlm2024 ; alayrac2022flamingo for convergence, making these impractical for resource and data limited settings. Moreover, gains from using multiple visual encoders often come at the cost of latency during inference. In this work, we hypothesize that scaling the number of visual inputs and/or training data improves visual perception performance owing to its positive effect on the visual representations inside the LLM. Therefore, we explore a hidden opportunity to optimize the visual quality of representations inside the LLM directly. To that end, we propose to distill knowledge from a set of expert visual encoders into the LLM’s representations during the PT stage through a set of embedding losses (Fig. 1c). During inference, we only use a single encoder, resulting in a better trade-off between visual understanding performance and efficiency as compared to feeding multiple visual inputs to the LLM (Fig. 1a,b).
We first conduct extensive experiments to establish a relationship between the (M)LLM’s visual representation quality and its downstream VQA performance. To that end, we train a probe at each layer inside the LLM (of the MLLM, LLaVA-1.5 liu2023improvedllava , in our case due to its wide usage in the academic community) to analyze their quality against expert visual features. We choose features from visual encoders trained for three tasks as targets: image segmentation jain2023oneformer ; sam , depth estimation depth_anything_v1 ; depth_anything_v2 , and image generation ldm ; unclip , owing to their well-studied, fundamental nature and the first two being seminal perception abilities he2017maskrcnn ; eigen2014depth . Based on our choice of target features, we refer to benchmarks tong2024cambrian1 evaluating the depth and spatial reasoning ability as our target tasks, i.e., we aim to improve the model at visual perception without harming its general reasoning abilities. We find that the LLM representations also improve with more data, indicating enhanced visual perception ability and downstream performance, proving the effectiveness of our probing setup (Fig. 2).
From our probing experiments, we also observe that the middle LLM layers are optimal for embedding visual information inside the LLM based on the layer-wise representation quality trend. Consequently, we investigate the effect of optimizing the intermediate LLM representations against the expert visual features at specific layers. Inspired by the recent works in embedding predictive architectures for self-supervised learning jepa , we propose VisPer-LM (Visual Perception Language Model) to distill kd knowledge from the expert visual encoders into the LLM’s representations during the pre-training (PT) stage. Specifically, we optimize an embedding loss between the target feature and the embedding predictor output at the corresponding LLM layer (Fig. 1c). Note that we still use features from a base encoder as inputs to the LLM. Furthermore, we incorporate a specialized set of tokens, $\langle t\rangle$, enriched with target information into the LLM’s input sequence, fostering an implicit visual chain of thought goyal2024think ; icot while enhancing the model’s ability to handle spatial and depth reasoning queries. Our experiments in Sec. 5 illustrate the effectiveness of our approach on various benchmarks while outperforming the baselines. To summarize, our contributions are:
  • Through probing existing MLLMs, we observe a positive correlation between visual representation quality and downstream performance on benchmarks like CV-Bench. To the best of our knowledge, we are the first to analyze the visual quality of MLLMs’ representations.
  • We present VisPer-LM, an approach to distill knowledge from expert visual encoders into LLM’s representations to improve the model’s visual (depth/spatial) perception abilities.
  • We conducted extensive experiments to demonstrate the superiority of VisPer-LM over the corresponding single and multi-encoder baselines on various benchmarks, including boosts up to 8. 7% and 5. 6% on the depth and distance task in CV-Bench, respectively.
2 Related Works
Figure 3: Probing Visual Representation Quality across LLM layers in MLLMs. (1) As shown in the first row, the multi-encoder baseline has the best probing performance owing to the additional feature inputs. The performance of probes trained on our VisPer-LM falls between the two baselines, demonstrating the effectiveness of our embedding distillation approach in learning an improved projector while only using a single encoder during inference. (2) We observe that the probing performance for single-encoder models trained solely with natural language supervision improves as the training data of the base MLLM increases, indicating that the LLM improves its visual representations of the world with more training data. In the last row, we observe that our VisPer-LM (base setting) outperforms LLaVA-1.5 trained with more data during the PT stage, demonstrating the effectiveness of our approach with limited (data/compute) resources.
2.1 MLLMs for Visual Reasoning
Contemporary MLLMs have three primary components: vision encoder(s), projector, and LLM. One line of work zhu2023minigpt ; llava-one ; liu2023improvedllava ; li2024cumo ; liu2023llava uses a single pretrained vision encoder clip ; vit ; convnext ; cherti2023reproducible ; siglip and trains a projector, like an MLP liu2023improvedllava or QFormer instructblip to align the visual features with the LLM. A few recent approaches chameleon ; diao2024EVE attempt to develop native MLLMs by directly feeding image patches into the LLM without using any visual encoder. Some works alayrac2022flamingo ; nvlm2024 train cross-attention modules inside the LLM, modifying the LLM architecture and often requiring millions of training data samples. Another line of work explores using either multiple encoder features tong2024cambrian1 ; mgllava ; shi2024eagle ; li2024mgm or multiple visual modalities jain2024vcoder ; liu2024prismer ; cheng2024spatialrgpt as inputs to the LLM for improved visual (spatial) reasoning performance. Our approach uses a single-base vision encoder while distilling information from expert visual encoders into the LLM’s representations.
2.2 Perception Probing in Foundation Models
OthelloGPT othello probed the features from a GPT-2 gpt2 trained on sequences from a board game, Othello, and found that the probes were able to learn the board state, indicating the ability of sequence models to learn spatial world representations. A recent work probes the features from foundational vision encoders banani2024probing3dawarenessvisual for 3D tasks. In our work, we probe the representation from the LLM layers against visual features from expert perception models. Furthermore, we establish a relationship between visual representation quality inside the LLM and downstream VQA performance.
2.3 Self-Supervised Learning
Distilling information kd from a target encoder into a source encoder is a well-established technique to improve the source encoder’s embeddings for a downstream task simclr ; chen2020exploring ; grill2020bootstrap . Recently, I-JEPA jepa proposed an embedding predictive architecture to improve representations inside a source encoder by comparing the target encoder features and mapped source encoder features with a trained predictor. A concurrent work, REPA yu2024repa improved DiTs Peebles2022DiT at image generation by distilling information from DINOv2 oquab2023dinov2 . Recently, a few works radio ; diva first distill information from target encoders into a single model and then leverage the resulting model as the vision encoder in an MLLM, requiring relatively longer and more training data for distillation. Unlike previous works, we distill information from multiple vision encoders into the LLM embedding space during the PT stage, resulting in a more vision-centric training approach without any extra training data.
3 Visually Probing LLM’s Embeddings
Figure 4: Architecture for VisPer-LM. During Pre-Training (PT), we optimize an embedding loss at specific layers for each target encoder: layers $d\in\mathbb{D}$, $s\in\mathbb{S}$, and $g\in\mathbb{G}$ for the depth, segmentation, and generation tasks, respectively. We use a resampler-based embedding predictor perceiver , denoted as $\mathbf{P}^{l}_{\{\text{task}\}}$ at each layer $l$, to output predictions. Each predictor takes in two inputs: a set of learnable queries ($\mathbf{Q}^{{\text{task}}}$) and the token sequence from layer $l$, with special tokens for other tasks omitted. The task tokens are derived from the corresponding embedding predictor’s learnable queries. During IFT, we train with only the next-token prediction objective while keeping the special tokens frozen to not affect their nature as we found it to perform empirically better in Sec. 5.
In this section, we systematically analyze the quality of LLM representations within the MLLM through probing with the expert visual features as the targets. We define high-quality representations as those that map easily to the target feature space, demonstrated by a high cosine similarity to the corresponding expert encoder’s features. We choose target encoders from models trained for three visual tasks: image segmentation, generation, and depth estimation, guided by the mentioned tasks’ fundamental and interpretable nature, i.e., visualizing the representations’ quality through the respective decoders. We use the encoder outputs from Swin-L OneFormer jain2023oneformer ; swin-T , DINOv2-L Depth Anything-v2 vit ; depth_anything_v2 , and CLIP-ViT-L unCLIP-SD-2.1 unclip ; clip as the segmentation ($\mathbf{E}^{\text{seg}}$), depth ($\mathbf{E}^{\text{depth}}$) and generation ($\mathbf{E}^{\text{gen}}$) as target probe features, respectively.
Probing Setup. We train a single-layer Perceiver Resampler perceiver as the probe head at every LLM layer for each of the three target features. We use a resampler probe head (similar to the Emb Predictor shown in Fig. 4) to accommodate the different sequence lengths of the target features and representations inside the LLM. For each LLM layer, we input learnable queries into the probe head while using the layer’s hidden states (for all tokens) as keys for cross-attention inside the resampler block. During probing, we set the learning rate as $1e^{-3}$ and minimize the smooth-L1-loss objective with a batch size of 256. We train the probes for two epochs on the 118k images from the COCO-train2017 coco set with the text query: “Describe the image in two lines.”. We find a positive correlation of 0.98 between the depth probing performance and CV-Bench accuracy in Tab. X, proving the effectiveness of our probing setup. We analyze the following LLaVA-1.5-based models with Llama-3-8b dubey2024llama3herdmodels as the LLM and CLIP-ConvNeXT-XXL cherti2023reproducible as the visual encoder:
  • Single-Encoder MLLM with features from $\mathbf{E}^{\text{base}}$ passed through an MLP into the LLM.
  • Multi-Encoder MLLM with features from $\mathbf{E}^{\text{base}}$, $\mathbf{E}^{\text{gen}}$, $\mathbf{E}^{\text{depth}}$, and $\mathbf{E}^{\text{seg}}$ concatenated in the feature dimension (in order) and passed through an MLP into the LLM.
  • VisPer-LM with features from $\mathbf{E}^{\text{base}}$ passed through an MLP into the LLM and information from $\mathbf{E}^{\text{gen}}$, $\mathbf{E}^{\text{depth}}$, and $\mathbf{E}^{\text{seg}}$ distilled into the LLM representations.
We compute the cosine similarity between the probe outputs and the corresponding target features over the 5k images from the COCO-val2017 coco set to get the probing performance for evaluation.
Layerwise Trend. Upon evaluating the probes in the single encoder baseline, we observe that the middle (12–24) layer probes show the best representation quality for the depth and seg probing tasks, with an upward trend in quality in the initial layers and a downward trend in the deeper layers, as shown in Fig. 3. We attribute the observed trend to the fact that the middle layer representations contribute the most to the (visual) reasoning (primarily spatial/depth) in LLMs kaduri2024_vision_of_vlms ; kaplan2024tokenswords ; skean2024layer_rep . It is an important finding when deciding the position of the embedding losses, since optimizing the middle layer representations should be optimal. Surprisingly, the probes trained to predict the features from $\mathbf{E}^{\text{gen}}$ have fairly high cosine similarity (greater than 0.7) for all layers. We attribute the mentioned phenomenon to the choice of $\mathbf{E}^{\text{gen}}$ that already has language-aligned features, unlike $\mathbf{E}^{\text{depth}}$ and $\mathbf{E}^{\text{seg}}$.
Visual Encoding Approach. As shown in the first row of Fig. 3, the probes for multi-encoder MLLM expectedly learn better representations than the probes for the single-encoder MLLM. Interestingly, for the former, the depth and seg probe performance remain at the same level till layer 20 and then follow a downward trend, indicating the possibility of the deeper layer representations becoming text-rich in deeper layers lin2024boosting . We do not observe a downward trend for gen probing, owing to the text-aligned target gen features. We observe that the VisPer-LM probes fall between the two baselines, serving as a good trade-off between efficiency and visual representation quality.
Training Data. We study the effect of scaling training data on the visual representation quality inside the single-encoder MLLM in the second row of Fig. 3. We analyze four models: (i) PT: model after the PT stage, (ii) PT + 50% IFT: model with complete PT and trained on 50% IFT data, (iii) PT + IFT: model with complete PT and IFT, and (iv) PT + VPT + IFT: model with an extra training stage on ALLaVA-Caption chen2024allava data, during which the whole model is trained. We observe that with an increase in the amount of training data for the probed model, the probes show a gradual improvement, indicating that the representations of the visual world inside MLLMs and, therefore, downstream performance improve with just natural language supervision on more data (Fig. 2a)!
For experimental completeness, we also report probing for the single-encoder LLaVA-1.5 trained with additional ALLaVA-Caption chen2024allava data during the PT stage. As illustrated in the third row of Fig. 3, while the inclusion of additional PT data improves LLaVA-1.5’s performance on CV-Bench, our VisPer-LM achieves superior results with lesser PT data, highlighting the effectiveness of our approach with limited data. Notably, there is a significant improvement in representation quality for the seg probing task with the use of ALLaVA-Caption during PT, which we attribute to the dataset’s detailed captions fostering enhanced semantic alignment chen2024allava . Please refer to Tab. I for more results.
4 Embedding Visual Information into LLM
The current trend in developing MLLMs for visual reasoning is gluing vision encoder(s) to a decoder LLM with a projector and training the model to minimize the cross-entropy loss for classification over the LLM’s vocabulary for next token prediction (NTP). In our attempt to add a vision perspective to training MLLMs while only using a single visual encoder during inference, we aim to optimize a predictive visual embedding objective for the intermediate LLM representations with the standard NTP objective during the PT stage. In other words, we distill auxiliary visual (task) information into the LLM’s intermediate representations rather than feeding the visual features into the LLM from the expert encoders. We hypothesize that such an approach leads to a better projector initialization (Tab. VIII) for the IFT stage. Our hypothesis is validated in Fig. 3, as probes for our VisPer-LM significantly outperform the single encoder baseline, especially in the initial layers. During IFT, we train the MLLM using the NTP objective, without the embedding predictors.
4.1 Multi-Encoder Feature Distillation
Given a set of target visual encoders, $\mathbf{E}^{\text{depth}}$, $\mathbf{E}^{\text{seg}}$, and $\mathbf{E}^{\text{gen}}$, we distill information from their feature outputs into the LLM’s representation space by minimizing embedding losses between predictor outputs and target features at certain layers. Since the token sequence inside the LLM has a different length as compared to the target representations, we use a single-layer Perceiver Resampler perceiver as the embedding predictor ($\mathbf{P}^{l}_{\text{\{task\}}}$) that takes as inputs: learnable latent queries ($\mathbf{Q}^{\text{\{task\}}}$) along with the embeddings from layer $l$ of the LLM and outputs a prediction, $\mathbf{p}^{l}$. We set the number of queries in $\mathbf{Q}^{\text{\{task\}}}$ such that it matches the number of tokens in the corresponding $\mathbf{E}^{\text{\{task\}}}$ features, i.e., $576$, 576576, and $1$, for $\mathbf{Q}^{depth}$, $\mathbf{Q}^{seg}$, and $\mathbf{Q}^{gen}$, respectively.
To amplify the effect of the contextual information about the target task (spatial/depth reasoning) inside the LLM chen2023minigptv2 ; goyal2024think , we use a set of special $\mathbf{N}$ task tokens, $\langle t\rangle$. We append $\langle t\rangle$ to the image tokens in the LLM’s input sequence. Specifically, we average pool $\mathbf{Q}^{depth}$ and $\mathbf{Q}^{seg}$ into $\mathbf{N}$ number of tokens to obtain the $\langle d\rangle$ and $\langle s\rangle$ tokens, respectively. Since the number of target tokens for generation features is only one, we average pool $\langle g\rangle$ to obtain $\mathbf{Q}^{gen}$. During PT, we only train the MLP projector, embedding predictors, and the special tokens $\langle t\rangle$. We extract the tokens corresponding to the system prompt, input image, the corresponding $\langle t\rangle$, and the text query from a layer’s output sequence as the input keys to the embedding predictor, as shown in Fig. 4.
4.2 Predictive Embedding Optimization
At the core of our approach is indirectly optimizing the projector during the PT stage by minimizing an embedding loss for each target representation at specific layers. As shown in Fig. 4, we feed the outputs from the $\mathbf{d}\in\mathbb{D}$, $\mathbf{s}\in\mathbb{S}$, and $\mathbf{g}\in\mathbb{G}$ layers of the LLM into the corresponding embedding predictor to obtain the predictions for computing an embedding loss ($\mathcal{L}_{\text{emb}}$), which is a weighted sum of Smooth-L1-Loss fast-rcnn and contrastive (InfoNCE) loss oord2018representation . We denote the sets of layers for embedding distillation from $\mathbf{E}^{\text{depth}}$, $\mathbf{E}^{\text{seg}}$, and $\mathbf{E}^{\text{gen}}$ as $\mathbb{D}$, $\mathbb{S}$, and $\mathbb{G}$, respectively. The final embedding loss for each target feature is a sum of losses over all corresponding layers. We compute the final loss during PT as the sum of the NTP objective and embedding losses, as shown in Eq. 4.
We denote $\mathbf{p}^{l}$ and $\mathbf{t}$ as the embedding predictor ($\mathbf{P}^{l}_{\{\text{task}\}}$) outputs at layer $l$ and the target task features, respectively. $\mathbf{p}^{l}_{i}$ denotes the $i^{\text{th}}$ element in a batch of embeddings with a global batch size of $B$ aggregated over all GPUs. We denote $\tau$ (initialized to $2.0$) as the learnable scaling factor for the contrastive loss. The weights for smooth-L1-and contrastive losses are $\lambda_{\text{sL1}}=1$ and $\lambda_{\text{contrastive}}=0.3$, respectively, at all selected layers. We select these values to ensure their magnitudes are comparable and their weighted sum aligns in scale with $\mathcal{L}_{\text{NTP}}$ at convergence to maintain training stability. We set $\lambda_{\text{depth}}$, $\lambda_{\text{seg}}$, and $\lambda_{\text{gen}}$ as 0.5.
$\mathcal{L}^{l}_{\text{sL1}}(\mathbf{p}^{l},\mathbf{t})=0.5\cdot(\mathbf{p}^{l}-\mathbf{t})^{2}\texttt{ if }|\mathbf{p}^{l}-\mathbf{t}|<1\texttt{ else }|\mathbf{p}^{l}-\mathbf{t}|-0.5$ (1)
$\mathcal{L}^{l}_{\text{contrastive}}=-\log\frac{\exp(\text{sim}(\mathbf{p}^{l}_{i},\mathbf{t}_{i})/\tau)}{\sum_{j=1}^{B}\exp(\text{sim}(\mathbf{p}^{l}_{i},\mathbf{t}_{j})/\tau)}$ (2)
$\mathcal{L}^{\mathbb{D/S/G}}_{\text{emb}}=\sum_{l\in\mathbb{D/S/G}}\left(\lambda_{\text{sL1}}\mathcal{L}^{l}_{\text{sL1}}+\lambda_{\text{contrastive}}\mathcal{L}^{l}_{\text{contrastive}}\right)$ (3)
$\mathcal{L}_{\text{PT}}=\mathcal{L}_{\text{NTP}}+\lambda_{\text{depth}}\mathcal{L}^{\mathbb{D}}_{\text{emb}}+\lambda_{\text{seg}}\mathcal{L}^{\mathbb{S}}_{\text{emb}}+\lambda_{\text{gen}}\mathcal{L}^{\mathbb{G}}_{\text{emb}}$ (4)
Other Architecture Details. As shown in Fig. 4, we use DINOv2-L oquab2023dinov2 from Depth Anything v2 depth_anything_v2 as $\mathbf{E}^{\text{depth}}$, Swin-L swin-T from OneFormer jain2023oneformer (trained on COCO-train2017 coco ) as $\mathbf{E}^{\text{seg}}$, and CLIP-ViT-L clip from unCLIP-SD-2.1 unclip as $\mathbf{E}^{\text{gen}}$. In the input sequence to the LLM, we append the gen, depth, and seg tokens in that order to the image tokens. Therefore, given an image-text pair, the input token arrangement is $\langle sys\rangle|\langle img\rangle|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\langle g\rangle}|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\langle d\rangle}|{\color[rgb]{0.75,0.0,1.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0.0,1.0}\langle s\rangle}|\langle txt\rangle$}, where $\langle sys\rangle$, $\langle img\rangle$, and $\langle txt\rangle$ denote the tokens corresponding to the system prompt, image, and text query, respectively. The target feature dimensions from $\mathbf{E}^{\text{gen}}$, $\mathbf{E}^{\text{depth}}$, and $\mathbf{E}^{\text{seg}}$ are $(1,1024)$, $(576,1024)$, and $(576,1536)$, respectively, corresponding to their final layer outputs.
5 Experiments
Table 1: Comparisons to Baselines. Our VisPer-LM outperforms the single and multi-encoder LLaVA-1.5 liu2023improvedllava by up to 2.5% and 0.9% on average across various benchmarks, respectively. The best numbers are set in bold for every base-encoder and decoder LLM combination.
CV-Bench General
[…]
Phi3-4k-mini
LLaVA-1.5 CLIP-ViT-L 52.4 67.2 75.2 56.3 36.5 57.1 56.7 57.3
VisPer-LM (ours) CLIP-ViT-L 52.4 68.7 76.0 56.7 36.0 58.0 56.4 57.7
LLaVA-1.5 CLIP-ConvNeXT-XXL 51.8 70.8 74.0 55.3 36.4 58.0 55.9 57.4
VisPer-LM (ours) CLIP-ConvNeXT-XXL 49.4 72.5 77.2 60.3 38.4 58.4 56.5 58.9
Llama3-8b
LLaVA-1.5 CLIP-ViT-L 50.4 73.3 64.9 48.7 38.8 57.8 56.9 55.1
LLaVA-1.5 (feat concat.) CLIP-ViT-L + $\mathbf{E}^{\text{depth}}$ + $\mathbf{E}^{\text{seg}}$ + $\mathbf{E}^{\text{gen}}$ 45.3 75.5 70.9 54.3 36.1 57.5 58.3 56.8
LLaVA-1.5 (token concat.) CLIP-ViT-L + $\mathbf{E}^{\text{depth}}$ + $\mathbf{E}^{\text{seg}}$ + $\mathbf{E}^{\text{gen}}$ 45.9 75.7 68.9 52.7 37.8 56.5 59.3 56.7
VisPer-LM (ours) CLIP-ViT-L 51.3 74.2 69.4 54.3 39.5 57.9 56.6 57.6
Cambrian-1 tong2024cambrian1
CLIP-ConvNeXT-XXL + CLIP-ViT-L +
SigLIP-SO-400M siglip + DINOv2-G oquab2023dinov2
57.2 65.0 63.2 50.1 34.5 53.2
LLaVA-1.5 RADIO-ViT-L radio 56.4 64.5 65.8 50.1 36.6 55.2 57.4 55.1
LLaVA-1.5 CLIP-ConvNeXT-XXL 54.1 62.8 69.5 49.8 37.4 57.5 56.3 55.3
VisPer-LM (ours) CLIP-ConvNeXT-XXL 57.4 71.5 66.8 52.8 38.5 55.0 59.0 57.3
Table 2: Results on Additional Benchmarks. Our VisPer-LM outperforms LLaVA-1.5 on classical benchmarks like POPE pope and GQA gqa showing its intact general reasoning ability.
[…]
LLaVA-1.5 RADIO-ViT-L radio 86.3 62.8 36.9 50.5 59.1
LLaVA-1.5 CLIP-ViT-L 85.9 63.5 38.6 50.6 59.7
VisPer-LM (ours) CLIP-ViT-L 86.4 63.7 38.7 54.0 60.7
Table 3: Scalability with VPT (more training data). VisPer-LM outperforms LLaVA-1.5 on average across different CV-Bench tasks. We use CLIP-ConvNeXT-XXL as the base encoder.
[…]
LLaVA-1.5 Phi3-4k-mini 49.7 70.0 72.6 58.7 61.8
VisPer-LM (ours) Phi3-4k-mini 53.7 72.0 73.1 58.5 63.4
LLaVA-1.5 Llama3-8b 56.3 76.8 73.1 50.3 63.3
VisPer-LM (ours) Llama3-8b 60.0 75.0 70.8 55.2 64.6
In this section, we first provide a comprehensive comparison of our method’s performance to that of the base MLLM, LLaVA-1.5 liu2023improvedllava across different base vision encoder and decoder LLM choices in Tab. 1. Next, we provide experimental results with a 2.5-stage training strategy, composed of an extra Visual Pre-Training (VPT) stage that involves training on the ALLaVA-Caption-663K chen2024allava dataset to demonstrate the scalability of our approach to more data. Lastly, we methodically study various design factors, including the optimal choice of layers for embedding additional visual information and the number and nature of special tokens ($\langle t\rangle$) through a series of ablations.
5.1 Implementation Details
Training. During the PT stage, we use the LLaVA-558K liu2023improvedllava dataset to train our model for an epoch with $lr$ of $1e^{-3}$. We only train the (MLP) projector, the embedding predictors, and the special tokens ($\langle t\rangle$). During the IFT stage, we use the LLaVA-665K liu2023improvedllava dataset and train the projector and LLM for one epoch with an $lr$ of $2e^{-5}$ with the vision encoder and $\langle t\rangle$ kept frozen. When using VPT, we leverage the ALLaVA-Caption-663K chen2024allava dataset to train the whole model (except $\langle t\rangle$) for one epoch with an $lr$ of $2e^{-5}$. Unless mentioned otherwise, we report results with models trained with PT and IFT stages. During the PT stage, we train our model with the NTP objective and the embedding prediction objectives. During the VPT and IFT stages, we solely use the NTP objective.
We train all our models on 16 AMD 192G-MI300X GPUs with a batch size of 256 during PT and 128 during IFT and VPT. We use CLIP-ViT-L clip and Llama3-8b dubey2024llama3herdmodels as the base vision encoder and decoder LLM unless mentioned otherwise. By default, we set $\mathbf{N}$, $\mathbb{D}$, $\mathbb{S}$, and $\mathbb{G}$ to 8, $\{8,20\}$, $\{10,18\}$, and $\{12,20\}$, respectively. For other hyperparameters, we follow LLaVA-1.5 liu2023improvedllava .
Evaluation. We primarily evaluate VisPer-LM for vision-centric abilities on CV-Bench tong2024cambrian1 and report results for all four tasks in CV-Bench: Count (2D), Relation (2D), Depth (3D), and Distance (3D). For ablations, we report the average accuracy over the 2D (count and relation) and 3D (depth and distance) task categories. We also evaluate our models’ general visual reasoning capabilities on the MMStar mmstar , RWQA xai_grok_2024 , and OK-VQA okvqa benchmarks. We select MMStar as our primary benchmark for general visual reasoning, as it addresses the well-known limitations mmstar of existing benchmarks MMBench ; lu2022learn ; yue2023mmmu ; li2023seed through careful filtering. Nonetheless, we report results on several classical general reasoning benchmarks gqa ; pope ; vizwiz ; yue2023mmmu in Tab. 2 for experimental completeness.
Table 4: Comparison on the Perception tasks using BLINK fu2024blink benchmark. Our VisPer-LM improves significantly over LLaVA-1.5 liu2023improvedllava on depth and relation reasoning, our target perception tasks.
Method Avg Spatial Relation Relative Depth Count
LLaVA-1.5 55.6 72.0 51.4 43.3
RADIO-L LLaVA-1.5 55.8 65.7 52.4 49.2
Cambrian-1 56.9 74.1 51.6 45.0
VisPer-LM (ours) 58.8 75.5 51.6 49.2
Table 5: Ablations on Layer sets for $\mathcal{L}_{\text{emb}}$. Setting $\mathbb{D}$$\{8,20\}$, $\mathbb{S}$$\{10,18\}$, and $\mathbb{G}$$\{12,20\}$ performs the best. Following the findings in Sec. 3, we only experiment with middle layers.
$\mathbb{D}$ $\mathbb{S}$ $\mathbb{G}$ CV-Bench2D CV-Bench3D MMStar Avg
{20} {18} {20} 57.6 60.8 38.8 52.4
{8;20} {10;18} {12;20} 58.6 64.2 39.5 54.1
{18;20} {18;20} {16;20} 55.8 59.5 40.8 52.0
{16;18;20} {16;18;20} {16;18;20} 56.8 61.3 37.0 51.7
Table 6: Training stages for $\mathcal{L}_{\text{emb}}$. Using the embedding losses only during PT is optimal.
PT IFT CV-Bench2D CV-Bench3D MMStar RWQA Avg
single-encoder 56.0 61.0 38.8 57.8 53.4
57.7 62.9 38.8 57.5 54.2
58.6 64.2 39.5 57.9 55.1
59.1 58.3 38.3 56.2 53.0
5.2 Main Results
As shown in Tab. 1, our VisPer-LM outperforms the single encoder baseline, i.e., LLaVA-1.5 liu2023improvedllava across different base encoder cherti2023reproducible ; clip and decoder LLM dubey2024llama3herdmodels ; abdin2024phi3technicalreporthighly combinations. Specifically, Llama3-8b dubey2024llama3herdmodels based VisPer-LM outperforms LLaVA-1.5 by 5.6% and 4.5% on the Distance and Relation task, respectively for the CLIP-ViT-L base encoder and by 8.7% on the Depth task for the CLIP-ConvNext-XXL base encoder. Furthermore, we compare our approach to the corresponding two variants of multi-encoder baselines: (i) feat concat.: features from all encoders are concatenated along the feature dimension and passed into a single MLP; and (ii) token concat.: features from all encoders are first passed through separate MLPs and then concatenated along with token dimension after average pooling the sequence outputs from $\mathbf{E}^{\text{depth}}$ and $\mathbf{E}^{\text{seg}}$ into eight tokens each. As shown in Tab. 1, our VisPer-LM outperforms the multi-encoder baselines on average, showing the effectiveness of our approach while using a single encoder during inference. Note that the multi-encoder baselines are better than VisPer-LM on the Depth task in CV-Bench owing to their superior depth representation quality as observed in Sec. 3.
We also compare our VisPer-LM with a RADIO-ViT-L radio and Llama3-8b-based LLaVA-1.5. Unlike our approach, which distills information directly into the LLM representations, RADIO distills knowledge from multiple expert encoders into a single student encoder. As shown in Tab. 1, our VisPer-LM outperforms the RADIO-ViT-L-based LLaVA-1.5, underscoring the effectiveness of our approach over encoder distillation for developing MLLMs. Note that we train the RADIO-ViT-L-based model using the same setting as the single encoder baseline.
Additionally, we train a Llama3-8b based Cambrian-1 tong2024cambrian1 model (with a batch size of 512, following Tong et al. tong2024cambrian1 ) using the LLaVA-558K (PT) plus LLaVA-665K (IFT) dataset mix. As shown in Tab. 1, the Cambrian-1 model not only underperforms as compared to our VisPer-LM but also LLaVA-1.5, underscoring the effectiveness of our approach with limited data, as the publicly released Cambrian-1 models were trained on millions of samples, requiring immense resources.
Additional Benchmarks. Although not the target tasks for our method, we report results on classical visual reasoning benchmarks like POPE pope , GQA gqa , MMMU yue2023mmmu , and VizWiz vizwiz in Tab. 2. VisPer-LM outperforms LLaVA-1.5, demonstrating its superior visual perception ability without any loss in general reasoning abilities.
Scalability with VPT. We report results using the 2.5-stage training setup (PT+VPT+IFT) for LLaVA-1.5 liu2023improvedllava and VisPer-LM in Tab. 3. Our VisPer-LM outperforms the baselines on average across all CV-Bench tasks, highlighting the scalability of our approach with more training data.
Results on BLINK fu2024blink benchmark. We report results on visual perception tasks under the BLINK benchmark (val set) in Tab. 4. We also report results using the data-matched Cambrian-1 and RADIO-ViT-L-based LLaVA-1.5 models. We find that our VisPer-LM significantly outperforms all other baselines on average, especially on the Spatial Relation task, demonstrating the effectiveness and generalization of our approach.
5.3 Ablations
Layer Sets for Embedding Losses. The choice of layers in $\mathbb{D}$, $\mathbb{S}$, and $\mathbb{G}$ for the corresponding embedding losses is a crucial design choice with a significant effect on the performance. Based on the findings about middle layers in Sec. 3, we ablate on embedding loss positions only in the middle layers of the LLM. As shown in Tab. 5, setting $\mathbb{D}$, $\mathbb{S}$, and $\mathbb{G}$ as $\{8,20\}$, $\{10,18\}$, and $\{12,20\}$, respectively, performs the best overall for the 32-layer Llama3-8b.
Training Stage for Embedding Optimization. In Table 6, we observe that adding $\mathcal{L}_{\text{emb}}$ during IFT results in worse performance than doing so only during the PT stage. We attribute this to the interference of vision-centric supervision with task-aligned natural language supervision during IFT. Additionally, the second row of Table 6 shows that using learnable tokens in the sequence goyal2024think without $\mathcal{L}_{\text{emb}}$ slightly improves performance over the baseline but remains inferior to VisPer-LM.
Table 7: Ablation on the nature of special tokens during IFT. Keeping $\langle t\rangle$ frozen during IFT aids in keeping their task-specific nature intact, resulting in better performance.
$\langle t\rangle$ during IFT CV-Bench2D CV-Bench3D MMStar RWQA Avg
frozen 58.6 64.2 39.5 57.9 55.1
learnable 56.9 56.1 39.0 57.3 52.3
Table 8: Throughput Analysis. Our VisPer-LM has superior inference throughout compared to that of the multi-encoder baseline and similar throughput to that of the single-encoder baseline with much better performance.
inference single-encoder multi-encoder VisPer-LM
throughput (samples/sec) 9.92 $\pm$ 0.03 5.32 $\pm$ 0.02 9.86 $\pm$ 0.01
Table 9: Experiments with SigLIP-VIT-SO400M siglip as the base vision encoder. Our VisPer-LM improves significantly over LLaVA-1.5 liu2023improvedllava on depth and relation reasoning, our target perception tasks.
Method CV-Bench Count Depth Relation
LLaVA-1.5 56.3 47.8 60.3 57.4
VisPer-LM (ours) 57.3 (+1.0) 48.7 (+0.9) 63.7 (+3.4) 66.3 (+8.9)
Nature of Special tokens during IFT. As shown in Tab. 7, keeping the special tokens ($\langle t\rangle$) frozen during IFT performs better as compared to making them learnable. We attribute frozen tokens performing better to the gradients from natural language supervision interfering with the vision-centric information stored in the special tokens during IFT if those are set as learnable parameters.
Throughput Analysis. We demonstrate the superior inference throughput of VisPer-LM over the multi-encoder baseline in Tab. 8. We record the throughput on a single NVIDIA 80G A100 GPU for a single forward pass on the CV-Bench evaluation set with a batch size of 1. We report the mean and standard deviation across 10 runs. We use the CLIP-ConvNeXT-XXL cherti2023reproducible and Llama3-8b dubey2024llama3herdmodels based models for the throughput analysis.
Experiments with SigLIP siglip . In Tab. 1, we achieved performance gains with CLIP-ConvNeXT-XXL, one of the strongest visual encoders, as also found in Cambrian-1 tong2024cambrian1 (ranked second overall in their Tab. 2). Here, we experiment with SigLIP-ViT-SO400M (ranked first in Cambrian-1 tong2024cambrian1 ) as our base visual encoder. As expected, we notice performance boosts on the perception tasks in CV-Bench in Tab. 9, which have a high positive correlation to the depth and seg representation quality (Appendix D), underscoring the effectiveness of our approach with various vision backbones.
6 Conclusion
In this work, we probed MLLMs and established a positive correlation between visual representation quality within the LLM and downstream performance. Building on these insights, we introduced VisPer-LM, the first approach to distill knowledge from target encoders into the LLM via predictive embedding optimization during the pre-training stage, complementing the next token prediction objective. We validated that our embedding optimization results in better vision-language alignment before the IFT stage. Through extensive experiments, we demonstrate VisPer-LM’s superiority to the corresponding baselines in terms of both representation quality and VQA performance.
Limitations and Future Work. With VisPer-LM, we enhanced the quality of representations inside the LLM with guidance from expert visual perception encoders. That said, our work mainly focuses on improving perceptual reasoning without any loss of general reasoning abilities. Incorporating more general-purpose teacher encoders, such as InternViT chen2024far ; chen2023internvl , offers a promising pathway to improve general reasoning abilities. Secondly, in this work, we mainly work with the image modality. Applying predictive embedding optimization for low-level information like motion control vjepa ; 2502.02492 while training on videos could improve MLLMs’ spatial and temporal reasoning in the future. Lastly, due to resource constraints, we could not scale VisPer-LM to larger LLMs like Llama3-70b dubey2024llama3herdmodels or Qwen2-72b qwen2 , and it remains an intriguing experiment.
Acknowledgements.
We extend our heartfelt gratitude to Fangrui Zhu, Reuben Tan, Min Shi, Bhavika Devnani, Fiona Ryan, and Chieh-Yun Chen for their valuable feedback and insightful discussions. We also sincerely thank the GCR team at Microsoft for their support in resolving frequent infrastructure challenges, which enabled our experimentation. This work was in part supported by NSF CAREER Award #2239840, and the National AI Institute for Exceptional Education (Award #2229873) by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education. Lastly, we thank the ML Center @Georgia Tech and Microsoft Research for supporting this work.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Please refer to Sec. 3 and Sec. 5 for verifying the claims.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Please refer to the end of Sec. 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not include any theoretical results or claims.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Please refer to Sec. 3 for details about our probing experiments and to Sec. 4 and Sec. 5.1 for our main experiments.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes] for data; [No] for code.
  24. Justification: We use publicly available datasets which can be easily downloaded from the web. Our codebase is built on top of the LLaVA-1.5 liu2023improvedllava codebase, and we will release our code with the camera-ready version.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Please refer to Sec. 5.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We conduct all our experiments with a fixed seed, resulting in a deterministic nature of our results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Please refer to Sec. 5.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: To the best of our knowledge, we follow the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We do not believe our work has any societal impacts except the usual hallucination concern involved with the usage of MLLMs. Our embedding distillation approach does not affect this in any form or way.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: We shall release the code and models under a non-commercial license upon acceptance to ensure only research use.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have credited all prior works with citations and will include the corresponding license in our code release upon acceptance.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release any new asset.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We did not use LLMs except for training LLM-based MLLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
In this appendix, we first share additional ablations, including the effect of the order of different special tokens ($\langle g\rangle$$\langle d\rangle$$\langle s\rangle$) in the input sequence to the LLM and the different key input possibilities to the embedding predictor in Appendix A. We use CLIP-ViT-L clip and Llama3-8b dubey2024llama3herdmodels as the base vision encoder and decoder LLM, respectively, for the ablations, unless mentioned otherwise. Next, we demonstrate the effectiveness of our probing setup for encoder-free MLLMs in Appendix C and establish a clear correlation between probing performance and CV-Bench accuracy in Appendix D. Lastly, we provide qualitative and quantitative analysis on the downstream probing tasks in Appendix E.
Appendix A Additional Ablations
Table I: Using additional data during PT v/s embedding optimization. Our VisPer-LM demonstrates superior performance than the LLaVA-1.5 model trained on with additional ALLaVA-Caption chen2024allava data during the PT stage, underscoring the effectiveness of our approach with limited data. We use CLIP-ConvNeXT-XXL cherti2023reproducible ; convnext as the base visual encoder.
Method PT IFT CV-Bench2D CV-Bench3D MMStar OK-VQA Avg
LLaVA-1.5 LLaVA-558K LLaVA-665k 60.0 56.3 37.4 56.0 52.4
LLaVA-1.5 LLaVA-558K + ALLaVA-Caption-663K LLaVA-665k 56.8 60.8 37.1 57.5 53.1
VisPer-LM LLaVA-558K LLaVA-665k 60.8 62.2 38.5 59.0 55.1
Table II: Setting $\mathbf{N}=8$ is optimal with $\mathbf{N}=0$ setting also giving boosts on CV-Bench and MMStar.
$\mathbf{N}$ CV-Bench2D CV-Bench3D MMStar RWQA Avg
single-encoder 56.0 61.0 38.8 57.8 53.4
0 56.1 62.0 40.1 56.3 53.6
8 58.6 64.2 39.5 57.9 55.1
16 56.6 63.6 37.1 54.5 52.9
24 55.7 60.0 39.3 57.4 53.1
Table III: Key input to the Embedding Predictor.. Feeding the tokens corresponding to the system prompt, image, corresponding special tokens, and the text query is optimal.
key input to emb. predictor CV-Bench2D CV-Bench3D MMStar RWQA Avg
$\langle img\rangle$ $|$ $\langle t\rangle$ 53.0 54.6 38.4 56.7 50.7
$\langle sys\rangle$ $|$ $\langle img\rangle$ $|$ $\langle t\rangle$ 58.7 63.0 38.8 57.4 54.5
$\langle sys\rangle$ $|$ $\langle img\rangle$ $|$ $\langle t\rangle$ $|$ $\langle txt\rangle$ 58.6 64.2 39.5 57.9 55.1
Number of Special Tokens. We analyze the effect of the number of special tokens per target feature ($\langle t\rangle$) on the performance in Tab. II. Setting $\mathbf{N}$ as eight results in the best average performance across benchmarks. Furthermore, VisPer-LM without $\langle t\rangle$ also outperforms the baseline on CV-Bench3D and MMStar, respectively, demonstrating the effectiveness of our embedding optimization.
Input tokens to Embedding Predictor. As shown in Tab. III, we find that including the tokens corresponding to the system prompt in the key input to the embedding predictor is critical for performance. We attribute it to system tokens having high attention scores and effect on the generation xiao2023streamingllm . Therefore, distilling target information into the system tokens is crucial for performance. Moreover, including the text query tokens in the key input to the embedding predictors also results in a slight performance boost as the text tokens hold global image information kaduri2024_vision_of_vlms .
Table IV: Embedding Optimization Modes. Using the depth, seg, and gen embedding losses simultaneously is optimal.
mode CV-Bench2D CV-Bench3D MMStar Avg
LLaVA-1.5 56.0 61.0 38.8 51.9
depth 58.6 63.5 38.8 53.6
seg 56.2 57.6 38.2 50.7
gen 56.2 65.8 39.3 53.8
depth + seg 58.6 61.8 38.6 53.0
depth + gen 53.6 61.8 38.8 51.4
seg + gen 54.2 60.2 39.3 51.2
depth + seg + gen 58.6 64.2 39.5 54.1
Table V: Order of different special tokens in the input sequence to the LLM. Appending the gen, depth, and seg tokens (in that order) in the LLM’s input sequence after the image tokens is the optimal setup.
[…]
LLaVA-1.5 50.4 73.3 64.9 48.7 58.5
$\langle d\rangle$ $|$ $\langle s\rangle$ $|$ $\langle g\rangle$ 49.4 68.7 69.2 56.2 59.9
$\langle d\rangle$ $|$ $\langle g\rangle$ $|$ $\langle s\rangle$ 51.6 72.8 70.3 54.5 61.4
$\langle s\rangle$ $|$ $\langle d\rangle$ $|$ $\langle g\rangle$ 48.7 71.3 65.2 52.5 58.5
$\langle s\rangle$ $|$ $\langle g\rangle$ $|$ $\langle d\rangle$ 46.7 71.3 71.2 50.8 58.9
$\langle g\rangle$ $|$ $\langle d\rangle$ $|$ $\langle s\rangle$ 51.3 74.2 69.4 54.3 61.4
$\langle g\rangle$ $|$ $\langle s\rangle$ $|$ $\langle d\rangle$ 50.9 68.8 70.0 50.5 59.2
Table VI: Embedding Loss weights during PT. Setting each embedding loss’ weight to 0.5 is optimal.
$\lambda_{\text{depth}}$ $\lambda_{\text{seg}}$ $\lambda_{\text{gen}}$ CV-Bench2D CV-Bench3D MMStar Avg
LLaVA-1.5 56.0 61.0 38.8 51.9
0.10 0.10 0.10 60.5 61.3 38.3 53.4
0.25 0.25 0.25 56.3 59.4 37.1 50.9
0.50 0.50 0.50 58.6 64.2 39.5 54.1
0.75 0.75 0.75 57.9 59.4 37.6 51.6
1.00 1.00 1.00 55.8 61.7 38.1 51.9
Table VII: Ablations on components of embedding losses. Using both smooth-L1-loss and contrastive loss to compute the final embedding loss is optimal.
$\mathcal{L}_{\text{sL1}}$ $\mathcal{L}_{\text{contrastive}}$ CV-Bench2D CV-Bench3D MMStar Avg
56.8 62.3 38.3 52.5
58.6 64.2 39.5 54.1
Embedding Optimization Mode. In this ablation study, we evaluate various combinations of embedding losses applied during pretraining (PT). Our results, summarized in Tab. IV, reveal that the optimal performance is achieved when all three embedding losses—depth, seg, and gen—are used together. Interestingly, we observe that utilizing only depth or gen embedding losses still leads to notable performance improvements, whereas relying solely on seg embedding loss does not yield significant gains. This suggests that different types of target information contribute uniquely to the distillation process. Investigating how the distillation of one type of target information influences the effectiveness of others presents an intriguing direction for future research.
Order of Special Tokens. In Tab. V, we ablate the order of different special tokens in the LLM’s input sequence. We find that $\langle g\rangle$ $|$ $\langle d\rangle$ $|$ $\langle s\rangle$}\} and $\langle d\rangle$ $|$ $\langle g\rangle$ $|$ $\langle s\rangle$}\} show the best performance on CV-Bench tong2024cambrian1 . We choose $\langle g\rangle$ $|$ $\langle d\rangle$ $|$ $\langle s\rangle$}\} as our default order due to its performance being better than the baseline on all sub-tasks in CV-Bench.
Embedding Loss weights. In Tab. VI, we ablate on different values of $\lambda_{\text{depth}}$, $\lambda_{\text{seg}}$, and $\lambda_{\text{gen}}$ during for the corresponding embedding losses during the pre-training stage. We find that setting each loss weight to 0.5 is optimal.
Effect of Contrastive Embedding Loss. In Tab. VII, analyze the impact of the contrastive loss component within the embedding loss. Our findings show that incorporating the contrastive loss significantly enhances performance, highlighting its positive influence on the model’s effectiveness. We keep the smooth L1 loss as a default component to ensure the embedding predictions maintain the same magnitude as the target features, which is crucial for meaningful visualization.
Table VIII: Enhanced Projector post Predictive Embedding Optimization. We observe that the projector features in VisPer-LM are better than those in LLaVA-1.5 after the PT stage (proj_PT), indicating a positive effect of the embedding losses during the PT stage on the learned projector. We use CLIP-ConvNeXT-XXL and Llama3-8B as the base visual encoder and decoder LLM, respectively.
Category Model Module Output Probed Stage Probe cosine-sim Scores
depth seg gen avg
vision_encoder CLIP-ConvNeXT-XXL 0.545 0.512 0.719 0.592
proj_PT LLaVA-1.5 Projector PT 0.531 0.495 0.781 0.602
VisPer-LM (ours) Projector PT 0.554 0.503 0.776 0.611
Qualitative Comparisons. We provide qualitative comparisons demonstrating the difference between LLaVA-1.5 liu2023improvedllava and VisPer-LM for the different tasks in CV-Bench in Fig. VI, Fig. VII, Fig. VIII, and Fig. IX.
Visualizing Embedding Predictions for Target Tasks. We visualize the visual quality of LLM representations after the PT stage for our VisPer-LM using the decoders from the corresponding target models, as shown in Fig. I. We observe that the decoder outputs have good object shape and boundary quality, demonstrating the successful representation optimization with our embedding losses.
Figure I: Visualizing Embedding Predictor Outputs after the PT stage. The quality of the decoded representations indicates the effectiveness of our embedding optimization.
Figure II: Ground-truth outputs from the target models used for Probing MLLMs.
Figure III: Probing Encoder-Free MLLMs. We probe Chameleon-7B chameleon and EVE-7B diao2024EVE models to show the effectiveness of our probing setup for encoder-free MLLMs. In the first row, we compare the mentioned models to LLaVA-1.5 and our VisPer-LM from Sec. 3 of the main text. On the one hand, we notice that EVE-7B shows worse probing performance in the initial layers than LLaVA-1.5, and the representation quality follows an upward trend even in deeper layers, resulting in features of similar quality to LLaVA-1.5, which is also reflected in the CV-Bench accuracy. On the other hand, Chameleon-7B has the best representation quality in the early middle layers, with the quality dropping significantly in the deeper layers. We include only the EVE-7B and Chameleon in the second row for better readability.
Table IX: Evaluating Encoder-Free MLLMs on CV-Bench. Our VisPer-LM significantly outperforms native encoder free VLMs like EVE-7B diao2024EVE and Chameleon-7B chameleon .
[…]
Chameleon chameleon Llama2-7b 13.6 51.7 49.5 50.8 39.6
EVE diao2024EVE Vicuna-7b 52.2 61.3 69.2 53.8 58.4
LLaVA-1.5 CLIP-ConvNeXT-XXL Llama3-8b 54.1 62.8 69.5 49.8 58.2
VisPer-LM (ours) CLIP-ConvNeXT-XXL Llama3-8b 57.4 71.5 66.8 52.8 61.5
Appendix B Enhanced Projector post Pre-Training
In this section, we probe the features from the visual encoder and the projector for the LLaVA-1.5 and VisPer-LM models from Sec. 3 of the main text, i.e., using CLIP-ConvNeXT-XXL and the Llama3-8B as the base visual encoder and decoder LLM, respectively. Note that only the projector is trained during the PT stage, while the projector and the LLM are trained during the IFT stage. The visual encoder is kept frozen during both stages.
We compare the probing scores for the feature outputs from the projector module of the LLaVA-1.5 and VisPer-LM models between checkpoints obtained after the PT stage (proj_PT) of training. As shown in Tab. VIII, the representations in models under proj_PT are better for VisPer-LM than LLaVA-1.5, indicating the predictive embedding optimization assisting in learning an improved projector before the IFT stage.
We observe that the quality of depth and seg representations in LLaVA-1.5 under proj_PT are worse than those of the base visual encoder (vision_encoder), indicating a loss of depth/spatial information after the features are fed into the projector in LLaVA-1.5. However, for VisPer-LM, the depth representations have better quality than those from the vision encoder. Note that OLA=VLM’s seg representations are slightly worse than those from CLIP-ConvNeXT-XXL but still better than those from LLaVA-1.5.
Appendix C Probing Encoder-Free MLLMs
Following the settings described in Sec. 3 of the main text, we probe all layers from the publicly available encoder-free Chameleon-7B chameleon and EVE-7B diao2024EVE model checkpoints against the depth, seg, and gen target features in Fig. III.
We notice that the representation quality for features from EVE-7B is relatively worse in the initial layers, owing to the patch embedding operation that outputs coarse features. However, the probing performance improves in the deeper layers, which could be attributed to the patch-alignment operation diao2024EVE . Notably, EVE-7B uses about 35M diao2024EVE training samples but still lags slightly behind LLaVA-1.5 in performance (Tab. IX), with our VisPer-LM showing better representation quality and performance, underscoring the effectiveness of our approach.
Unlike EVE-7B, features from Chameleon-7B chameleon have surprisingly low depth and seg representation quality in the deeper layers. Unlike other models, Chameleon-7B follows a steep downward trend in the deeper layers, which we attribute to its relatively low training on VQA data with a heavy focus on text-to-text and text-to-image tasks in the SFT data mixture, which is also evident from its low performance on CV-Bench. Note that during evaluation, we noticed that Chameleon cannot output option letters required to evaluate on MCQ-type CV-Bench. Therefore, we leverage Qwen-2.5-72B-Instruct qwen2 to extract the option letters corresponding to the sentence output from Chameleon to obtain the numbers for Tab. IX.
For both the models, gen probes perform relatively better due to the text-aligned nature of the gen target features.
These findings provide critical insights into developing future encoder-free MLLMs, re-establishing the importance of VQA data in the SFT mixture, and supervising the vision tokens inside the LLM from an expert encoder during training. Our proposed embedding optimization opens up a promising avenue to this end.
Appendix D Correlation between VQA Performance and Probing Performance
Takeaways 1. As shown in Tab. X, we observe the highest positive correlation between CV-Bench accuracy and seg/depth probing performance. 2. Moreover, we find a high correlation between the performance on the Depth/Relation task and Count tasks with the depth and seg probing performance, respectively. 3. As expected, we find low performance correlation with gen probe scores, indicating that gen features act as a good proxy for general visual representation targets. 4. Surprisingly, we observe a negative correlation between the Distance task and all probe scores, which could be attributed to the lack of “real-world distance” estimation queries in the training dataset.
Table X: Correlation of CV-Bench performance to probing performance. We notice the high positive correlation of 0.98 between depth probing performance and CV-Bench accuracy. Moreover, we find that performance on the Count and Depth subtasks also highly correlates with the seg and depth probe cosine-similarity scores, respectively. Additionally, we find a low correlation between the gen probing performance and CV-Bench performance.
[…]
depth 0.90 0.98 0.70 -0.31 0.96 0.94 0.98
seg 0.98 0.93 0.50 -0.15 0.91 0.96 0.98
gen 0.34 0.46 0.27 -0.31 0.37 0.36 0.37
Figure IV: Plotting Correlation between Probing performance and CV-Bench accuracy. We find a high correlation between the eval cosine-sim scores and the accuracies on the CV-Bench as well as the 2D and 3D categories under CV-Bench for the depth and seg probes, validating their choice as the target task to improve the MLLM’s perception ability.
Figure V: Plotting Correlation between Probing performance and CV-Bench sub-tasks accuracy. We find a high correlation between the depth eval cosine-sim scores and the Depth/Relation/Count task performances. We make the same observation about seg probe scores and the Count/Depth task. Interestingly, we find a low negative correlation between the probe scores and Distance task performance.
Figure VI: Qualitative Examples for the Count task in CV-Bench. Our VisPer-LM can accurately predict the presence of one picture and one tree, unlike LLaVA-1.5 liu2023improvedllava .
In this section, we analyze the correlation between the performance on CV-Bench (evaluates models’ visual perception ability) and probing performance (measure of model’s representation quality), basing our conclusion on trends for six models and their corresponding depth/seg/gen probes:
We follow the same settings as mentioned in Sec. 3 and Sec. 5 of the main text to train our probes and models, respectively.
Probing Mode: depth. We observe a high correlation of 0.98 between the CV-Bench accuracy and depth probe eval cosine-sim scores, as shown in Fig. IV. It is critical proof of our claim that improved visual representations and the corresponding VQA performance go hand-in-hand. We also find a correlation with a 0.98 and 0.80 factor between the accuracy on the Depth and Relation tasks (in CV-bench), respectively, with the depth probe scores in Fig. V, explaining the gains achieved by VisPer-LM.
Probing Mode: seg. We also observe a high correlation of 0.98 between the CV-Bench accuracy and seg probe eval cosine-sim scores, as shown in Fig. IV. The high correlation indicates that CV-Bench requires strong perception abilities, validating our decision to adopt it as the primary benchmark for our work. We also find a correlation of 0.98 between the accuracy on the Count task (in CV-bench) with the seg probe scores in Fig. V.
Probing Mode: gen. We observe that the gen probe scores do not correlate highly with CV-Bench or any of its tasks, as shown in Fig. IV and Fig. V. We attribute this finding to the generative features’ text-aligned nature, which is not optimal for developing the model’s perception abilities.
Appendix E Evaluating Probes on Downstream Tasks
We evaluate the probes trained against the target features on the corresponding downstream target tasks, i.e., image generation, depth estimation and image segmentation. To obtain the predictions, we feed the outputs from the probes into the decoder from the corresponding target model. Specifically, we report the FID Seitzer2020FID scores on 5k images from COCO-val2017 coco for gen probes, accuracy on the DA-2K depth_anything_v2 benchmark for depth probes, and mIoU jain2023oneformer on the COCO-val2017 coco set for seg probes. We average the scores over all layers for easier comparison. As shown in Tab. XI, probes trained for our VisPer-LM outperform those for the baseline LLaVA-1.5 liu2023improvedllava model across all probing tasks, proving the improved visual representation quality owing to our embedding optimization approach.
Figure VII: Qualitative Examples for the Depth task in CV-Bench. Our VisPer-LM can accurately predict that the lamp and keyboard ar closer to the camera in the respective samples.
Figure VIII: Qualitative Examples for the Relation task in CV-Bench. Our VisPer-LM can accurately predict that the positions of the trees and the bottle in the respective samples.
Figure IX: Qualitative Examples for the Distance task in CV-Bench. Our VisPer-LM can accurately predict that the distances between the respective pair of objects.
Table XI: Quantitative Evaluation on target probing task. Probes trained for our VisPer-LM perform significantly better as compared to the probes trained on baseline LLaVA-1.5 liu2023improvedllava .
Probed Model FID Seitzer2020FID ($\downarrow$) DA-2K % Acc. depth_anything_v2 ($\uparrow$) % mIoU jain2023oneformer ($\uparrow$)
LLaVA-1.5 23.1 66.4 39.3
VisPer-LM (ours) 22.4 77.8 45.4
Target Encoder 18.1 97.3 64.5
Figure X: Layerwise visualizations for the depth probes. For LLaVA-1.5 liu2023improvedllava , the probes generate blob-like outputs up to the eighth layer, with visualizations progressively improving in the middle layers, aligning with the findings presented in Sec. 3 of the main text. Notably, probes trained on VisPer-LM begin producing distinguishable object shapes and boundaries as early as the third layer, attributed to the incorporation of embedding losses leading to improved representations in the initial layers.
Visualizing Probe Outputs. We present visualizations for the outputs for the probes trained on the single-encoder LLaVA-1.5 model in Sec. 3 in the main text. We provide the ground-truth visualizations from the teacher models in Fig. II.
As shown in Fig. X, the probe visualizations for the first eight layers of LLaVA-1.5 exhibit blob-like patterns, while the later layers progressively enhance the shape and boundary details of the foreground objects. In contrast, the probe visualizations for our VisPer-LM demonstrate improved object shapes and boundaries starting as early as layer-4, consistent with the findings on representation quality and layer-wise trends discussed in Sec. 3 of the main text. Additionally, we present probe visualizations for the seg and gen representations in Fig. XI and Fig. XII, respectively.
Figure XI: Layerwise visualizations for the seg probes. The LLaVA-1.5 probes often fail to segment the third car in the background for the first sample during the initial layers (layers two to eight), whereas the VisPer-LM probes demonstrate relatively better performance in this scenario.
Figure XII: Layerwise visualizations for the gen probes. The probe outputs for both the models are of fairly good quality.